OpenStudy (anonymous):

what is 3^(-2)-2? and 3^(-1)-2

OpenStudy (anonymous):

\[\frac{ 1 }{ 3^{2} }-2\] \[\frac{ 1 }{ 9 }-\frac{ 18 }{ 9 }=\frac{ -17 }{ 9 }\]

OpenStudy (anonymous):

remember the exponential rule \[a^{-m}={1\over a^m}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Since its raised to a negative it should become: \[\frac{ 1 }{ 3^2 }-2\]

OpenStudy (anonymous):

a negative exponent is called a "reciprocal"

OpenStudy (anonymous):

you can figure out the second one using the rule electrokid just mentioned

OpenStudy (anonymous):

yes thank you @electrokid

OpenStudy (anonymous):

Yeah @electrokid gave you a great 'definition' from which you can work thorough the next problem.

OpenStudy (anonymous):

thanks